OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (anonymous):

Sorry I didn't mean to my laptop messed up.

hero (hero):

j = t + 10 j - 3 = 2(t-3) j = t + 10 j = 2t - 3 j = j t + 10 = 2t - 3 2t - t = 10 + 3 t = 13 j = 23 Three years ago, t = 10, j = 20

OpenStudy (anonymous):

How old is tom?

OpenStudy (anonymous):

now

hero (hero):

Tom = 13